Didadas manufactures, among other things, running shoes. Their size 8 men’s shoes are the most in demand, and it is important that these are all exactly size 8. The cost for deviation can be determined by the standard quadratic quality loss function using k = $30. The current production process generates pairs of shoes with an average size of 8.2 and a standard deviation of 0.1. If each production run consists of 50,000 pairs of shoes, then what is the total expected loss for each production run?

a) $75,000

b) $225,000

c) $1.50

d) $17,000
Answer #1

1. A. 75000

D = PROCESS MEAN - SPECIFICATION = 8.2 - 8 = 0.199999999999999
EL(x) = K * (STANDARD DEVIATION^2 + D^2) = 30 * (0.1^2 + 0.199999999999999^2) = 1.5

FOR 50000 UNITS:

50000 * 1.5 = 75000

** Leaving a thumbs-up would really help me out. Let me know if you face any problems.